The composition of air is approximately 75.52% N_2
Answer #1

(a) dS = - R (x1 ln(x1) + x2 ln(x2) + x3ln(x3) + x4ln(x4)

dS = - 8.314 J/mol K * (0.7552*ln(0.7552) + 0.2315*ln(0.2315) + 0.0128*ln(0.0128) + 0.00046*ln(0.00046))

dS = 5.07 J/mol K

(b) Assume we have a mixture of 100 mol of air. If we Ar and CO2 frozen, we lost:

1.28 + 0.046 moles = 1.326 moles

And we have now:

100 - 1.326 = 98.674 moles of air.

In this case, the new molar compositions are:

XN2 = 75.52 moles / 98.674 = 0.765

XO2 = 1- XN2 = 0.235

And the new entropy is:

dS = - R (x1 ln(x1) + x2 ln(x2)) = -8.314 J/mol L * (0.765*ln(0.765) + 0.235*ln(0.235))

dS = 4.53 J/mol K